A flask with a volume of 4.25 L contains 12.0 grams of an unknown gas at 25.0 C and 1.00 atm. What is the molar mass of the gas. R = 0.0821
1
Expert's answer
2021-10-08T02:08:42-0400

PV=nRT

1×4.25= n×0.0821×298

= 4.25= n×24.47

4.25= 24.47n

n= 0.174

Molar mass = 12/0.174

= 68.97g/mol